OpenStudy (anonymous):

The interest that Julia earns on her investment is given by the equation I = 48t where t is the time in years that her money is invested. She makes a table to determine how much money she would earn for different values of t. She selects 5 as a value for t. What is the corresponding table entry for I?

OpenStudy (bibby):

what is 48t where t is 5?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

oh oh nvm uhm 240

OpenStudy (bibby):

whenever 2 terms are stuck together without a sign in between it's usually multiplication. \(\huge 2y=2*y=2(y)=2 \times y\) correct

OpenStudy (anonymous):

Ah thanks bibsu

OpenStudy (bibby):

is that my pet name

OpenStudy (anonymous):

Yes

OpenStudy (bibby):

thank

OpenStudy (bibby):

a single thank to you

OpenStudy (anonymous):

welc